# Automated Statistical Model Discovery with Language Models

### v1

@@ Abstract
Statistical model discovery involves a challenging search over a vast space of models subject to domain-specific modeling constraints. Efficiently searching over this space requires human expertise in modeling and the problem domain. Motivated by the domain knowledge and programming capabilities of large language models (LMs), we introduce a method for language model driven automated statistical model discovery. We cast our automated procedure within the framework of Box’s Loop: the LM iterates between proposing statistical models represented as probabilistic programs, acting as a modeler, and critiquing those models, acting as a domain expert. By leveraging LMs, we do not have to define a domain-specific language of models or design a handcrafted search procedure, key restrictions of previous systems. We evaluate our method in three common settings in probabilistic modeling: searching within a restricted space of models, searching over an open-ended space, and improving classic models under natural language constraints (e.g., this model should be interpretable to an ecologist). Our method matches the performance of previous systems, identifies models on par with human expert designed models, and extends classic models in interpretable ways. Our results highlight the promise of LM driven model discovery.
@@ 1 Introduction
Figure 1: Language model driven automated model discovery. 1) The prompt for the LM contains the dataset in textual (or visual) form, (optionally) dataset metadata (e.g., dataset description), the code for previous probabilistic programs, and natural language feedback. 2) Given this, the LM proposes new models expressed as probabilistic programs. 3) We fit these programs using automated inference procedures provided by a probabilistic programming language. 4) After we fit models, we compute model criticism statistics that capture limitations of a modeling approach. We provide these statistics to a LM which produces natural language feedback to guide future proposals. 5) We propagate the best programs, their model criticism statistics, and natural language feedback, by inserting this information into the prompt.
Automated model discovery is not a new ambition. Previous systems have been successfully deployed for discovering physical laws (Bongard & Lipson, 2007; McKinney et al., 2006; Linka et al., 2023), reverse-engineering non-linear dynamical systems (Schmidt & Lipson, 2009), nonparametric regression (Duvenaud et al., 2013) and unsupervised learning (Grosse, 2014). However, in these systems, a human expert had to carefully design a domain specific language (DSL) of models and specify a hand-crafted search procedure for composing models in that DSL. For example, in the Automated Statistician (Duvenaud et al., 2013; Lloyd et al., 2014), a system for time series modeling, human experts defined a DSL of Gaussian process kernels and a search procedure that composes kernels via addition and multiplication. Defining the DSL and the operators for composing models requires significant modeling expertise. These systems also compromised flexibility for automation: rather than choosing a model class best-suited for a problem, experts chose models that compose conveniently. This breaks the core principle of separating modeling from inference (van de Meent et al., 2021).
@@
Leveraging LMs for automated model discovery is enticing at a conceptual level. However, in order to deliver on this promising idea, we need to make several important design choices. These design choices should exploit the strengths of LMs but remain grounded in principled statistical modeling. First, we need a representation of a statistical model that can be expressed programmatically and is amenable to automated inference (e.g., model fitting). Second, we need a method for guiding LM proposals through natural language.
Algorithm 1 Automated Model Discovery with LLMs
0:  dataset $\mathcal{D}$, number of rounds $T$, $k$ number of exemplars, $m$ number of proposals per round, (optional) warm-start example $z_{0}$, function for scoring a program score, (optional) function for producing natural language feedback synthesize
@@
     $\{z_{t}\}_{i=1}^{m}\sim q_{\text{LM}}(\cdot|\mathcal{Z},z_{0},h^{t-1},\mathcal{D})$
     $\{s_{t}\}_{i=1}^{m}\leftarrow\texttt{score-all}(\texttt{score},\{z_{t}\}_{i=1}^{m},\mathcal{D})$
     $\mathcal{Z}\leftarrow\texttt{select-exemplars}(k,\{z_{t}\}_{i=1}^{k},\{s_{t}\}_{i=1}^{m})$
     $h_{t}\leftarrow\texttt{synthesize}(k,\{z_{t}\}_{i=1}^{k},\{s_{t}\}_{i=1}^{k})$
@@
We evaluate our method in three settings that cover common use cases in probabilistic modeling: searching within a DSL, searching over an open-ended space of probabilistic models, and improving expert models subject to modeling constraints expressed in natural language. In the first use case, we illustrate that our LM system is effective even in a DSL and matches the performance of the Automated Statistician (Duvenaud et al., 2013). We then consider the more general setting of automatically constructing probabilistic models for real world data; crucially, we do not require a user to define a DSL and this generality is enabled by our choice to use probabilistic programs. Our method identifies probabilistic programs that match the performance of human expert written programs. In the third setting, we use LMs to improve classic models and illustrate a compelling advantage of using LMs for model discovery: we can express modeling constraints in natural language (e.g., this model should be more physical) that guide LMs towards models that balance interpretability and flexibility.
2 Box’s Loop with Language Models
@@ 2.1 Background
Before introducing our method, we provide a brief overview of the probabilistic modeling paradigm. We then motivate a key design choice: representing probabilistic models with probabilistic programs. In probabilistic modeling, our goal is to describe a dataset in terms of unobserved, latent structure. Formally, a probabilistic model is a joint distribution $p(x,z|\eta);$ where $x=x_{1:N}$ denotes $N$ observed data points, $z=z_{1:M}$ denotes $M$ latent variables, and $\eta$ corresponds to non-random quantities in the model (Blei, 2014). After specifying a probabilistic model, we fit the model to observed data. Fitting a model involves inference, or computing the posterior distribution $p(z|x)$. In the model criticism step, we evaluate the model by interrogating the posterior. In this work, we use a common model criticism technique known as a posterior predictive check: to identify discrepancies, we sample from the posterior predictive distribution and compare these samples against observed data (Gelman et al., 2013). Model building and model criticism typically take place over multiple iterations in a process known as Box’s Loop (Box & Hunter, 1962).
There are many different representations of a probabilistic model. Since our goal is automated model discovery, we use probabilistic programs. Probabilistic programming languages provide ways to flexibly represent probabilistic models as programs and support generic inference methods for any arbitrary program (Wood et al., 2014; Goodman et al., 2008; van de Meent et al., 2021). In the context of automated model discovery, these are highly desirable properties, since LMs can write code reliably and other representations of probabilistic models can require custom inference methods. Our approach of using LMs to generate probabilistic programs is related to the approach taken by Wong et al. (2023) but is motivated by a different problem. Our emphasis is on using LMs as a tool for developing statistical models of real-world datasets, while their focus was on integrating symbolic methods with LMs to enhance their reasoning capabilities.
@@ 2.2 Problem formulation
We formally describe our framework. Our framework is motivated by recent work in inductive reasoning with LMs (Qiu et al., 2024; Wang et al., 2024), integrating tools with LMs (Gao et al., 2023), and driving LMs via linguistic feedback (Shinn et al., 2023).
At a high level, we consider a method for learning probabilistic models from data that involves two steps: a proposal step and a criticism step [^1] To avoid overloading the word “model”, we will refer to language models as LMs.. Crucially, by learning a model, we mean searching over the space of model structures and not just learning the parameters of some fixed model class. In each step, we leverage LMs. In the proposal step, a proposal LM proposes probabilistic programs for a dataset. We then fit these probabilistic programs and evaluate them. In the criticism step, we provide the critic LLM with programs and criticism statistics and ask the critic LLM to provide insights that guide the next round of proposals.
We start with a dataset $\mathcal{D}=\{\mathbf{x_{i}},y_{i}\}_{i=1}^{n}$. Here $\mathbf{x_{i}}\in\mathbb{R}^{d}$ are fixed $d$-dimensional input values (e.g., features) and $y_{i}\in\mathbb{R}$ are the observations. Let $\Sigma$ be the vocabulary of the LM. For each dataset, we have an associated metadata set $\mathcal{C}\in\Sigma^{d+1}$, which consists of natural language descriptions of $\mathcal{D}$ (e.g., animal ages vs length) and natural language descriptions of each feature in $\mathcal{D}$ (e.g., length of animal). Context informs how human modelers leverage prior knowledge; for example, if a modeler knows their dataset consists of monthly carbon dioxide measurements over a fifty-year time span, they will choose a model that can capture periodicity and linearity. Our goal is to find a probabilistic program $z\sim\Sigma^{*}$ that maximizes some notion of quality. We quantify this with the log marginal likelihood or expected log predictive density (ELPD) estimated via cross validation (LOO) (Vehtari et al., 2017).
@@ 2.3 Approach
To generate candidate models for round $t$, we sample $m$ probabilistic programs $z^{t}_{i}$ from the proposal LM $q_{\text{LM}}(\cdot)$. In our experiments, we use GPT-4 V (Achiam et al., 2023), which has multimodal capabilities. We leverage in-context learning, or LMs ability to learn from examples in a prompt, to guide the LM’s proposals based on high-scoring programs in the past (Brown et al., 2020). The proposal LM $q_{\text{LM}}$ “conditions” on $h^{t-1}\in\Sigma^{*}$, a natural language instruction synthesizing previous modeling approaches and suggesting new approaches, $k$ exemplars $\{z_{1},\ldots z_{k}\}$, and a visual or textual representation of $\mathcal{D}$. In later analyses, the proposal LM also conditions on a warm-start expert program $z_{0}$.
$z_{i}^{t}\sim q_{\text{LM}}(\cdot|z_{0},z_{1},\ldots,z_{k},h^{t-1},\mathcal{D})$
Chain-of-thought reasoning, or generating intermediate reasoning steps, improves the performance of LMs (Wei et al., 2022; Kojima et al., 2022). Motivated by this, we instruct the proposal LM to reflect on the properties of the dataset in the plot of the data, sketch a high-level modeling approach, state the hypotheses that it will address before writing a program, and add comments to code that address specific hypotheses.
To create the round $t$ exemplars for the Proposal LM, we choose the best $k$ programs among the $m$ proposed programs in round $t-1$. Alternatively, we could choose the $k$ best programs across all rounds. However, we sometimes noticed that this alternative approach suffers from a pathology: all future proposals might collapse to a single “ancestor” program from an early round. This can happen if a better model class is proposed in round $t$ (e.g., hierarchical model) but the specific parameterization under-performs a model proposed earlier (e.g., non-hierarchical).
Scoring Step
In the scoring step, we fit a probabilistic program to data. Fitting this model requires approximate inference so we leverage pymc (Abril-Pla et al., 2023), a probabilistic programming library in Python. pymc automatically assigns a Markov Chain Monte Carlo (MCMC) sampler to perform inference. Crucially, by using a probabilistic program, we decouple modeling from inference: the proposal LM focuses on constructing the best model, while pymc takes care of inference. Our approach of offloading computations to an external tool is connected to recent work augmenting LM capabilities through interaction with tools (e.g., Python interpreter) (Gao et al., 2023; Schick et al., 2023).
@@
In the criticism step, we ask the LM to identify weaknesses in the models proposed so far. To help the LM identify weaknesses, we compute model criticism statistics for a list of previously proposed probabilistic programs. We then provide these model criticism statistics to a critic LM to identify what modeling approaches worked. Naively, we could simply provide all proposed programs to the critic LM. However, this list grows each round and potentially has redundancy. Furthermore, LMs struggle to reason over long contexts. On the other hand, only providing the top programs in a given round might overly bias the LM to the most recent round of proposals. Since probabilistic programs can sometimes be summarized succinctly in natural language (e.g., von Bertalanffy growth function with informative priors), we distill information about previous approaches by asking a critic LM to provide linguistic feedback $h^{t-1}$ on programs (Shinn et al., 2023). In particular, to produce $h^{t-1}$, we sample from a critic LLM $p_{LM}$ and again leverage in-context learning. Let $\tilde{z}_{1},\ldots\tilde{z}_{d}$ be the top $d$ programs across all rounds of revision. We generate $h^{t-1}$ as follows
$h^{t-1}\sim p_{\text{LM}}(\cdot|\tilde{z_{1}},\ldots,\tilde{z_{k}},h,\mathcal{D})$
Here, $h$ is an instruction to the LM for what kind of feedback to provide. In our experiments, we ask the LM to 1) provide natural language hypotheses for discrepancies between the model and the data and 2) provide a synthesis of what strategies/approaches worked. Crucially, to capture long-term dependencies and reduce bias towards the recent round, the programs $\{\tilde{z_{i}}\}_{i=1}^{d}$ provided to the critic LM are the best $d$ programs across all rounds; in general, this need not overlap with the exemplars $\{z_{j}\}_{j=1}^{m}$ provided to the proposal LM.

### v2

@@ Abstract
Statistical model discovery is a challenging search over a vast space of models subject to domain-specific constraints. Efficiently searching over this space requires expertise in modeling and the problem domain. Motivated by the domain knowledge and programming capabilities of large language models (LMs), we introduce a method for language model driven automated statistical model discovery. We cast our automated procedure within the principled framework of Box’s Loop: the LM iterates between proposing statistical models represented as probabilistic programs, acting as a modeler, and critiquing those models, acting as a domain expert. By leveraging LMs, we do not have to define a domain-specific language of models or design a handcrafted search procedure, which are key restrictions of previous systems. We evaluate our method in three settings in probabilistic modeling: searching within a restricted space of models, searching over an open-ended space, and improving expert models under natural language constraints (e.g., this model should be interpretable to an ecologist). Our method identifies models on par with human expert designed models and extends classic models in interpretable ways. Our results highlight the promise of LM-driven model discovery.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Language model driven automated model discovery (BoxLM). 1) The prompt for the LM contains the dataset in visual and/or textual form, dataset metadata (e.g., dataset description), the code for previous probabilistic programs, and natural language feedback. 2) Given this, the proposal LM proposes new models expressed as probabilistic programs. 3) To fit these programs in a generic way, we leverage probabilistic programming languages and obtain scores and posterior predictive samples. 4) After we fit models, we compute the posterior predictive mean and variance. We provide these statistics to a critic LM which produces natural language feedback to guide the next round of model building. 5) We propagate the best programs, their posterior predictive means and variances, and natural language feedback forward by updating the prompt.
Automated model discovery is not a new ambition. Previous systems have been successfully deployed for discovering physical laws (Bongard & Lipson, 2007; McKinney et al., 2006; Linka et al., 2023), reverse-engineering non-linear dynamical systems (Schmidt & Lipson, 2009), nonparametric regression (Duvenaud et al., 2013) and unsupervised learning (Grosse, 2014). However, in these systems, a human expert had to carefully design a domain specific language (DSL) of models and specify a hand-crafted search procedure for composing models in that DSL. For example, in the Automatic Statistician (Duvenaud et al., 2013; Lloyd et al., 2014), a system for nonparametric regression and time series modeling, human experts defined a DSL of Gaussian process kernels and a search procedure that composes kernels via addition and multiplication. Defining the DSL and the operators for composing models requires significant modeling expertise. These systems also compromised flexibility for automation: rather than choosing a model class best-suited for a problem, experts chose models that compose conveniently. This breaks the core principle of separation of modeling from inference (van de Meent et al., 2021).
@@
Leveraging LMs for automated model discovery is enticing at a conceptual level. However, in order to deliver on this promising idea, we need to make several important design choices. These design choices should exploit the strengths of LMs but remain grounded in principled statistical modeling. First, we need a generic and flexible representation of a statistical model that can be expressed programmatically and is amenable to automated inference (e.g., model fitting). Second, we need a method for guiding LM proposals through natural language.
Algorithm 1 Automated Model Discovery with LMs
0:  dataset $\mathcal{D}$, number of rounds $T$, $k$ number of exemplars, $m$ number of proposals per round, (optional) warm-start example $z_{0}$, function for scoring a program score, (optional) function for producing natural language feedback criticize
@@
     $\{z^{t}_{i}\}_{i=1}^{m}\sim q_{\text{LM}}(\cdot|\mathcal{Z},z_{0},h^{t},\mathcal{D})$
     $\{s_{i}\}_{i=1}^{m}\leftarrow\texttt{score-all}(\texttt{score},\{z^{t}_{i}\}_{i=1}^{m},\mathcal{D})$
     $\mathcal{Z}\leftarrow\texttt{select-exemplars}(k,\{z^{t}_{i}\}_{i=1}^{m},\{s_{i}\}_{i=1}^{m})$
     $h^{t+1}\leftarrow\texttt{criticize}(\{z^{t}_{i}\}_{i=1}^{m},\{s_{i}\}_{i=1}^{m},h^{t})$
@@
We evaluate our method in three settings that cover common use cases in probabilistic modeling: searching within a DSL, searching over an open-ended space of probabilistic models, and improving expert models subject to modeling constraints expressed in natural language. In the first use case, we illustrate that our LM system is effective even in a DSL and matches the performance of the Automatic Statistician (Duvenaud et al., 2013). We then consider the more general setting of automatically constructing probabilistic models for real world data; crucially, we do not require a user to define a DSL and this generality is enabled by our choice to use probabilistic programs. Our method identifies probabilistic programs that match the performance of human expert written programs. In the third setting, we use LMs to improve classic models and illustrate a compelling advantage of using LMs for model discovery: given that certain modeling constraints can be difficult to express formally but easy to express in natural language (e.g., this model should be more physical), we use natural language to guide LMs towards models that balance interpretability and flexibility.
2 Automated Box’s Loop with Language Models
We begin with a brief background on the probabilistic modeling paradigm. We then formally introduce our problem setting and describe our approach. For an overview, see Figure 1 and Algorithm 7.
@@ 2.1 Background
In probabilistic modeling, our goal is to describe a dataset in terms of unobserved, latent structure. We describe a dataset through a probabilistic model which is a joint distribution $p(x,z|\eta)$; here $x=x_{1:N}$ denotes $N$ observed data points, $z=z_{1:M}$ denotes $M$ latent variables, and $\eta$ corresponds to non-random quantities in the model (Blei, 2014). After specifying a probabilistic model, we fit the model to observed data. Fitting a model involves inference, or conditioning on observed data and computing the posterior distribution $p(z|x)$. After fitting a model, we perform model criticism. In the model criticism step, we evaluate the model by interrogating the posterior. In this work, our model criticism is inspired by a common model criticism technique known as a posterior predictive check: to identify discrepancies, samples are drawn from the posterior predictive distribution and statistics of these samples are compared against those of the observed data (Gelman et al., 2013). Model building and model criticism typically take place over multiple iterations in an iterative process known as Box’s Loop (Box & Hunter, 1962).
There are many different representations of a probabilistic model. Since our goal is automated model discovery, we use probabilistic programs. Probabilistic programming languages provide ways to flexibly represent probabilistic models as programs and support generic inference methods for any arbitrary program (Wood et al., 2014; Goodman et al., 2008; van de Meent et al., 2021). In the context of automated model discovery, these are highly desirable properties, since LMs can write code reliably and other representations of probabilistic models can require custom inference methods. Our approach of using LMs to generate probabilistic programs is related to the approach taken by Wong et al. (2023) but is motivated by a different problem. Our emphasis is on using LMs as a tool for developing statistical models of real-world datasets, while their focus was on integrating symbolic methods with LMs.
@@ 2.2 Problem formulation
Our framework is motivated by recent work in inductive reasoning with LMs (Qiu et al., 2024; Wang et al., 2024), integrating tools with LMs (Gao et al., 2023), and driving LMs via linguistic feedback (Shinn et al., 2023).
At a high level, we consider a method for learning probabilistic models from data that involves two steps: a model building step and a criticism step [^1] To avoid overloading the word “model”, we will refer to language models as LMs.. Crucially, by learning a model, we mean searching over a space of model structures and not just learning the parameters of some fixed model class. In each step, we leverage LMs. In the proposal step, a proposal LM proposes probabilistic programs for a dataset. We then fit these probabilistic programs and evaluate them. In the criticism step, we provide a critic LM with programs and statistics assessing model fit (e.g., model criticism statistics) and ask the critic LM to provide feedback to guide the next round of proposals.
We start with a dataset $\mathcal{D}=\{\mathbf{x_{i}},y_{i}\}_{i=1}^{n}$. Here $\mathbf{x_{i}}\in\mathbb{R}^{d}$ are fixed $d$-dimensional input values (e.g., features) and $y_{i}\in\mathbb{R}$ are the observations. Let $\Sigma$ be the vocabulary of the LM. For each dataset, we have an associated metadata set $\mathcal{C}\in\Sigma^{d+1}$, which consists of natural language descriptions of $\mathcal{D}$ (e.g., animal ages vs length) and natural language descriptions of each feature in $\mathcal{D}$ (e.g., length of animal). Context informs how human modelers leverage prior knowledge; for example, if a modeler knows their dataset consists of monthly carbon dioxide measurements over a fifty-year time span, they will choose a model that can capture periodicity and a linear trend. Our goal is to find a probabilistic program $z\sim\Sigma^{*}$ that maximizes some notion of quality, which we take here to be either the log marginal likelihood or expected log predictive density (ELPD) estimated via cross validation (LOO) (Vehtari et al., 2017).
@@ 2.3 Approach
In the model building step, we automatically generate probabilistic programs for modeling a dataset given information about the dataset and previously proposed programs. In particular, to generate candidates for round $t$, we sample $m$ probabilistic programs $z^{t}_{i}$ from the proposal LM, $q_{\text{LM}}(\cdot)$. In our experiments, we use GPT-4 V (Achiam et al., 2023) (gpt4-11-06-preview), which has multimodal capabilities. We leverage in-context learning, or LM’s ability to learn from examples in a prompt, to guide the LM’s proposals based on high-scoring programs in the past (Brown et al., 2020). Specifically, $q_{\text{LM}}$ “conditions” on $h^{t}\in\Sigma^{*}$, a natural language instruction synthesizing previous modeling approaches and suggesting new approaches, $k$ exemplars $\{z_{1},\ldots z_{k}\}$, and a visual or textual representation of $\mathcal{D}$. Optionally, $q_{\text{LM}}$ also conditions on a warm-start expert program $z_{0}$:
$z_{i}^{t}\sim q_{\text{LM}}(\cdot|z_{0},z_{1},\ldots,z_{k},h^{t-1},\mathcal{D}).$
We run this at a temperature of 0.7. Chain-of-thought reasoning, or generating intermediate reasoning steps, improves the performance of LMs (Wei et al., 2022; Kojima et al., 2022). Motivated by this, we instruct $q_{\text{LM}}$ to reflect on the properties of the dataset or plot of the data, sketch a high-level modeling approach, state the hypotheses that it will address before writing a program, and add comments to code that address specific hypotheses.
To create exemplars $z_{1},\dots,z_{k}$ for round $t$ for $q_{\text{LM}}$, we choose the best $k$ programs among the $m$ proposed programs in round $t-1$.
Model Fitting Step
In the model fitting step, we fit a probabilistic program to data. This requires us to perform (approximate) inference for generic probabilistic models. To accomplish this, we leverage pymc (Abril-Pla et al., 2023), a Python probabilistic programming library. pymc automatically assigns a Markov Chain Monte Carlo (MCMC) sampler to perform inference; by default pymc uses a Hamiltonian Monte Carlo sampler (Homan & Gelman, 2014). Crucially, by using a probabilistic program, we decouple modeling from inference: the proposal LM’s role is to build a good model, while pymc takes care of inference. Our approach of offloading computations to an external tool is connected to recent work enhancing LMs by giving them tools (e.g., Python interpreter) (Gao et al., 2023; Schick et al., 2023).
@@
In the criticism step, we ask the critic LM, $p_{LM}$, to produce natural language criticism of fitted models; we use this criticism to drive model revision. First, we can obtain a scalar score measuring the model fit (e.g., ELPD LOO) for each proposed model. Second, to enable the critic LM to do something akin to a posterior predictive check, we obtain samples from the posterior predictive distribution (e.g., $p(x^{\prime}|z,x)=\int_{z}p(x^{\prime}|z)p(z|x)$) and then compute summary statistics of these posterior predictive samples (Gelman et al., 2013). For simplicity, we compute the posterior predictive means and variances for each fitted probabilistic program. We then provide $p_{LM}$ with select probabilistic programs, their scores $\mathcal{S}$ (e.g., ELPD LOO), the posterior predictive means and variances $\mathcal{P}$, and the dataset itself $\mathcal{D}$; we explore both visual and textual dataframe based representations of the posterior predictive and the dataset. Finally, we ask $p_{LM}$ to distill this criticism in natural language (e.g., von Bertalanffy growth function with informative priors) which we use to drive the next round of model building (Shinn et al., 2023).
The key design choice is which programs to provide to $p_{LM}$ (e.g., critic exemplars). Naively, we could provide all proposed programs across all rounds to $p_{LM}$. However, this list grows each round and has redundancy. Furthermore, LMs struggle to reason over long contexts. We therefore explore a simple approach to selecting exemplars where we provide $p_{LM}$ with the top $d$ programs $\tilde{z}_{1},\ldots\tilde{z}_{d}$ from the current round $t$. To produce $h^{t+1}$, we sample:
$h^{t+1}\sim p_{\text{LM}}(\cdot|\tilde{z}_{1},\ldots,\tilde{z}_{d},\mathcal{D},\mathcal{S},\mathcal{P}).$
In the appendix, we report additional results for an approach inspired by a state-space update (Baum & Petrie, 1966; Kalman, 1960; Rabiner, 1989). To avoid storing the entire history of proposed programs, we interpret $h^{t+1}$ as a latent state and compute it using the previous state $h^{t}$ and the new fitted programs $\{z_{i}\}_{i=1}^{m}$ at round $t$:
$h^{t+1}\sim p_{\text{LM}}(\cdot|{z_{1}},\ldots,{z_{m}},h^{t},\mathcal{D},\mathcal{S},\mathcal{P}).$
In practice, we implement this by asking $p_{\text{LM}}$ to add and delete hypotheses. We find that these two approaches have similar performances. We run this criticism step at a temperature of 0.0 to limit stochasticity in the criticism produced.
We refer to our full LM-driven automated Box’s loop as BoxLM.